OpenStudy (anonymous):

Solve 2e^5t = 75 plz help

OpenStudy (mrnood):

take logs (use natural logs here) note ln e^n = n

OpenStudy (anonymous):

2e5t=75 e5t=75/2 5t=ln(75/2) t=ln(75/2)/5 t=0.725

OpenStudy (anonymous):

Oh thank u

OpenStudy (mrnood):

@Rohith15 - Maybe giving the full answer is not truly helpful? Giving the method allows the OP to gain some understanding of the process....